OpenStudy (anonymous):

A cylinder labeled X has a height of 18 inches and a radius of 21 inches. Another cylinder labeled Y has a height of 90 inches and a radius of 105 inches. The volume of cylinder Y is _______ times the volume of cylinder X.

OpenStudy (anonymous):

Is the answer 125?

OpenStudy (anonymous):

\[V_2=kV_1\] \[2\pi (105^2) (90)=k(2\pi(21^2)(18))\] \[(21^2)(18)=k(105^2)(90)\] \[992250=7938k\] \[\frac{992250}{7938}=k\] 125

OpenStudy (anonymous):

Thanks!